# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v3

@@ IV-B2 Perception Information Selection
TABLE V: Summary of the Works of Temporal and Spatial Alignment.
ACD FFNet [185] 2023 V2I LiDAR DAIR-V2X Intermediate Detection $\surd$
SyncNet [186] 2022 V2V LiDAR V2V-Sim Intermediate Detection $\surd$
@@
APE RANSAC-based [189] 2022 V2V LiDAR COMAP Late Detection, Location $\times$
RobustV2VNet [190] 2021 V2V LiDAR V2V-Sim Intermediate Detection, Motion Prediction $\times$
@@
CBM [195] 2023 V2V Camera,LiDAR SIND[196], OPV2V Late Detection $\surd$
@@
ACD & APE V2X-ViT [198] 2022 V2X LiDAR V2XSet Intermediate Detection $\surd$
  • Note: ACD: Alignment for Communication Delay, APE: Alignment for Pose Errors.
IV-B3 Joint Agent and Information Selection
Methods that consider both cooperative agent and information selection can offer a more balanced approach to optimizing precision and bandwidth in CP. For example, Where2comm [172] introduced a spatial-confidence-aware CP framework that utilizes spatial confidence maps generated for each agent to identify perceptually critical areas. These critical features are then compactly packaged as messages and shared with others via a sparsely connected communication graph. A communication graph construction strategy is employed to determine with whom to communicate to minimize unnecessary bandwidth usage. After communication, confidence-aware multi-head attention is applied to enhance each agent’s feature map by incorporating received messages from other agents. Experimental results demonstrate that Where2comm efficiently enables mutually beneficial cooperation by sharing spatially sparse yet critical features in V2X communication.
IV-C Cooperative Information Alignment
To enable precise information fusion in CP, it is crucial to temporally and spatially align the perceptual information received via V2X communication with the ego data [199]. While many existing methods aim to synchronize exchanged messages within a consistent spatial coordinate system to address this challenge, practical issues such as misalignment and inconsistency can arise due to communication delays and pose noise [200]. Therefore, mitigating the effects of latency and pose errors is essential rather than relying solely on direct synchronization of asynchronous information. In this subsection, we provide a detailed discussion of all the aspects outlined in Table V.
IV-C1 Alignment for Communication Delays
In practical V2X communication scenarios, latency issues are inevitable due to various real-world factors, including limitations in wireless communication, model complexity, and the number of agents involved [201]. The presence of communication latency can mislead detectors, resulting in significant deviations in perception results and ultimately causing fusion misalignment. Temporal asynchrony can detrimentally affect the performance of the CP system, potentially leading to outcomes worse than those of single-vehicle perception [202]. Hence, developing a robust CP system capable of handling unavoidable communication delays is imperative.
A common and effective strategy to address the latency issue is to predict the sender’s perception at the current time step and subsequently perform coordinate transformations using the predicted data frame. In this context, the FFNet approach [185] exploits the temporal coherence of sequential infrastructure frames to predict future features and counteract asynchrony. FFNet adopts a similar architectural framework to that of SECOND [203], which includes a Backbone and Feature Pyramid Network, to extract the first-order derivative of feature flow from historical infrastructure frames. However, it is important to note that this first-order expansion method is most effective when the time interval is very short.
An alternative approach, proposed in [186], introduces a latency compensation module called SyncNet. SyncNet utilizes time series prediction techniques to estimate received features at the current time step based on historical information. To achieve this, a dual-branch pyramid Long Short-term Memory (LSTM) network is employed to infer real-time features and corresponding cooperative attention. Additionally, a time modulation module is integrated to allocate temporal attention between the estimated cooperative feature and the received asynchronous features, taking into account the latency time. The final estimation is generated by combining the estimated feature with the latest asynchronous feature based on the temporal weight.
Another approach for mitigating the impact of communication delays involves re-encoding received features with delay time. For instance, V2X-ViT [198] addresses time delay through a Delay-aware Positional Encoding (DPE) module that utilizes sinusoid functions conditioned on delay time and channel information to initialize a learnable embedding. This embedding is further updated through a linear projection. The resulting projected embedding, serving as an extension of temporal delay, is integrated into the received feature, thus achieving temporal alignment of features. In the case of V2VNet [187], each vehicle employs a fully connected GNN as an aggregation module, where each node represents the state of a CAV. Within this GNN, a CNN is utilized to compensate for latency, taking inputs such as the received intermediate feature, relative 6DoF pose, and delay time. The outcome is a time-delay-compensated representation concatenated with zeros to enhance the node’s ability to aggregate information from other vehicles.
Previous works have not addressed the issues raised by common irregular time delays, rendering existing collaborative perception systems unable to reach their full potential in real-world scenarios. As a substitute, CoBEVFlow [188] considered each collaboration message as one irregular sample and used irregular time series forecasting to estimate the BEV flow map for asynchronous feature alignment. In this way, the historical frames were used to estimate a BEV flow map, which encodes the motion information in each grid cell. The asynchronous perceptual features can be reassigned to appropriate spatial locations based on the BEV flow map and region of interest generation, thereby aligning perceptual features on the time dimension and mitigating the impact caused by asynchrony.
IV-C2 Alignment for Pose Errors
Establishing precise estimates of relative transformation matrices is foundational for V2X cooperation, ensuring consistent alignment of received perception information in a shared spatial coordinate system. In practical scenarios, these transformation matrices rely on sensor measurements from posture detection systems, such as the Global Positioning System (GPS), Real-Time Kinematic (RTK), and Inertial Measurement Unit (IMU). However, real-world systems are prone to imperfections, including noise and limited accuracy, which can introduce errors into the transformation process, potentially degrading perception performance.
Efforts have been made to address these challenges using statistical learning models. For example, one approach [189] corrected relative locations between two CAVs by matching key points in their perception results with the Random Sample Consensus (RANSAC) algorithm [204]. This method used the consensus of anchor points, such as vehicle centres and static object points, to calculate relative pose transformations accurately. While effective, this approach required the setting of problem-specific thresholds.
Within the framework of V2VNet [187], a consistency module [190] was introduced to improve relative pose estimations. This module structured the global coherence of absolute poses as a Markov random field, incorporating Bayesian reweighting. However, it is important to note that this model relies on ground-truth poses during the training phase, which may be limited in practical scenarios. In contrast, CoAlign [191] proposed an Agent-Object Pose Graph Optimization method for enhancing pose consistency without requiring precise supervision during training. This approach utilized an agent-object pose graph to model relationships and applied graph optimization techniques to achieve accurate pose alignment. Notably, CoAlign assumed that pose noise followed a Gaussian distribution, a simplifying assumption to facilitate the process.
The Iterative Closest Point (ICP) [205] algorithm is widely used for aligning 3D point clouds. In [192], ICP was employed to align two sets of points, where one set represented 3D bounding boxes detected by the Ego, and the other set represented those detected by the sender. This source and target point alignment were achieved using an ICP-based matching algorithm, iteratively optimizing the pose transformation to achieve a locally optimal alignment. However, it is important to note that ICP can be sensitive to points lacking correspondences, and its neighbour selection step relies on a parameter that directly impacts the registration outcome. In response to this challenge, the relative transformation was addressed as an optimal transport problem [206], seeking to minimize transportation costs between the source and target points. Nevertheless, the effectiveness of these approaches is contingent on the number of targets present within the shared FoV of both the Ego and the sender.
@@ IV-C2 Alignment for Pose Errors
High-definition (HD) maps represent a valuable resource for achieving precise spatial alignment, primarily due to their ability to provide accurate self-vehicle localization. In this regard, a map-based CP framework called the “map container” was introduced in [193]. This framework utilizes the HD map to establish precise benchmarks for spatiotemporal transformations of perceptual data. The core concept involves automatically transforming multi-agent perceptual information into the map coordinate space through a map-matching algorithm. This integration process uses diverse information sources within a decentralized fusion framework. The map container directly incorporates received information into the joint state estimation optimization process, showcasing a robust fusion architecture.
Utilizing multi-scale feature interaction can effectively mitigate feature map misalignment arising from pose errors. This approach involves encoding the feature map into multiple scales and subsequently aggregating them to derive the ultimate collaboration feature. To address this concern, a spatiotemporal awareness CP framework known as SCOPE was introduced in [194]. SCOPE aims to obtain robust representations while accounting for localization errors, achieved through confidence-aware multi-scale feature interaction. Specifically, the framework employs the spatial-confidence-aware method [172] to select pertinent features and generate confidence maps. These confidence maps are further encoded into three distinct scales. A pivotal component of SCOPE is the application of a deformable cross-attention layer [207], which facilitates the aggregation of information from key points extracted from the confidence maps. This aggregation process is governed by a threshold-based selection function applied at each scale. The final collaboration feature is generated by amalgamating information derived from these three scales through the utilization of a convolutional layer.
In the context of addressing the impact of GPS localization inaccuracies, the Multi-scale Window Attention Module (MSWin) introduced in V2X-ViT [198] plays a crucial role. MSWin is designed to learn long-range spatial interactions at various scales through a pyramid of resolution windows. Each of these windows accommodates a different attention range, enabling the simultaneous capture of both local and global spatial feature interactions. This approach enhances detection robustness in the face of positioning errors.
TABLE VI: Summary of the Works of Cooperative Information Fusion.
@@
Coop3D-Infra [110] 2022 V2I LiDAR CARLA Hybrid Detection $\surd$
@@
Recently, an innovative approach called inter-agent object association has emerged as a solution to mitigate localization errors in pose transformation. This method involves associating key points or objects within observations provided by connected agents, determining optimal relative pose transformation based on matched correspondences. In one approach presented in [195], a Context-based Matching (CBM) algorithm was introduced for inter-agent object association. CBM involves generating object clusters, considering each bounding box’s heading direction in local coordinate systems. Object correspondences are established through an iterative pairwise comparison procedure applied to all clusters to maximize global consensus. However, identifying correspondences can be challenging due to visual folding, non-visible objects, and noisy perception. To overcome these challenges, another method proposed in [197] introduces a deep masked graph-matching approach that frames correspondence identification as a graph-matching problem. This approach employs a masked GNN and includes a mechanism for handling non-visible objects. The network integrates multi-modal visual, spatial, and GPS information using SoftMax variance thresholding. Each node in the graph represents an object, while edges capture the spatial relationships between pairs of objects. This deep masked neural network explicitly addresses challenges related to object non-visibility caused by occlusion and limited field of view, ultimately enhancing the precision of correspondence identification.
IV-D Cooperative Information Fusion
@@
The effectiveness of CP crucially depends on the design of an efficient fusion strategy, which aims to consolidate the shared information to provide a coherent and accurate understanding of traffic situations. The following will examine recent fusion methods employed in CP information. For reference, Table VI provides a summary of all discussed fusion models in this subsection.
IV-D1 Traditional Fusion
Primary reduction operators like summation, maximum, or average pooling have found extensive use in cooperative information fusion for multi-agent perception. For instance, in [111], a raw-data CP system, dubbed Cooper system, was introduced for CAVs to enhance 3D object detection. However, this early cooperative approach entailed a substantial data overhead for transmitting raw point cloud data. F-Cooper [208] extended Cooper to the feature level to balance bandwidth constraints and detection accuracy in cooperative 3D object detection. Deep features were acquired through neural networks at each vehicle’s side, and max pooling was employed to combine V2V voxel features. Another model, PillarGrid [209], operated at the feature level for cooperative 3D object detection based on point cloud data. It employed a grid-wise feature fusion approach using max-pooling to integrate deep features from multiple agents. CoCa3D [210] leveraged point-wise maximum fusion to amalgamate BEV features from several CAVs to enhance camera-based detection through cooperative depth estimation. These CP methods outperformed single-vehicle detection techniques regarding 3D object detection performance.
In [211], a rule-based fusion scheme for CPMs was proposed. It identified conditional branches for associating locally detected objects with those detected via V2X communication. This fusion scheme was designed for scenarios where CAVs generated CPMs following standardized ETSI rules and transmitted them to a fusion centre via wireless communication. The fusion centre applied the proposed algorithm to generate merged CPMs, which were subsequently broadcast to all CAVs in range for downstream perception tasks. Simulation results at an intelligent junction illustrated the relevance and effectiveness of this fusion strategy for obstacle detection.
@@ IV-D1 Traditional Fusion
Non-parametric max operators and rule-based CPM fusion can integrate information with minimal computational cost but may fail to identify data correlations. Parametric methods, on the other hand, often yield optimal estimates by mathematically incorporating measurements and their associated uncertainties, resulting in more accurate and reliable fusion outcomes. For instance, the Covariance Intersection (CI) algorithm can consistently fuse estimates without prior knowledge of cross-correlations between them. Building upon this concept, a CI-based fusion strategy for CP was introduced [212]. This approach achieves a consistent fusion of local perception sensor data and perceived object information from other agents through V2X communication. However, it is important to note that high-level fusion methods, like the one described above, can be bandwidth-efficient. Still, they may encounter challenges when dealing with potentially noisy and incomplete single-agent perceptions, leading to less than satisfactory fusion results.
IV-D2 Graph-based Fusion
@@
Multi-agent collaboration can be effectively represented as a collaboration graph, where nodes represent individual agent states, and edges signify pairwise collaboration between agents. At the core of this collaboration graph lies the GNN, a vital component capable of extracting relational data representations and generating valuable node embeddings. These embeddings are derived from local node attributes and encompass information from neighbouring nodes within the graph. Consequently, node states can be continuously updated through inter-node communication and a neural network aggregation mechanism that combines received information.
In V2VNet [187], each vehicle employed a fully connected GNN [217] as an aggregation module, enabling intelligent data fusion from various viewpoints. Specifically, received messages were aggregated at each node using a mask-aware accumulation operator, ensuring consideration only of overlapping FoVs. Subsequently, a Convolutional Gated Recurrent Unit (ConvGRU) was utilized to update the state of each node and generate the fused intermediate representation. The gating mechanism of ConvGRU facilitated information selection for the accumulated received messages based on the current belief of the receiving CAV. The study showed that V2VNet successfully balanced accuracy improvements and bandwidth requirements. However, it employed a scalar-valued edge weight to indicate the importance of the neural message from one agent to another, which could not effectively model the strength of collaboration between two agents.
@@
On the other hand, DiscoNet [213] employed an edge encoder to establish a correlation between the ego feature map and the feature map from another agent, generating a matrix-valued edge weight. The edge encoder concatenated two feature maps along the channel dimension and applied four convolutional layers for downsampling. Additionally, a softmax operation was performed at each cell in the feature map to normalize the edge weights across multiple agents. Furthermore, knowledge distillation was utilized to enhance training by constraining post-cooperation feature maps in the intermediate fusion-based student model to align with the early fusion-based teacher model’s correspondences. This approach promoted improved feature abstraction and aggregation, enhancing object detection performance.
IV-D3 Attention-based Fusion
The attention mechanism introduced in the transformer model has found widespread use in determining fusion weights. Building on the concept of the heterogeneous graph transformer [218], the study presented in [198] introduced a V2X-based Vision Transformer (V2X-ViT). This model incorporates two key modules: the Heterogeneous Multi-agent Self-attention (HMSA) module and the Multi-scale Window (MSwin) attention module. The HMSA module discerns unique characteristics within the sensor measurements obtained from infrastructure and CAVs. In the HMSA module, agent types are associated with nodes and edges within a directed graph. Importance weights between pairs of nodes are computed based on the corresponding node and edge types. The MSwin module employs multiple window sizes to aggregate spatial information, thereby enhancing detection robustness in the presence of localization errors. Several V2X-ViT blocks are stacked to iteratively learn inter-agent interactions and per-agent spatial attention, ultimately yielding an aggregated feature representation for detection tasks. To further enhance object detection performance, HYDRO-3D [214] extends the V2X-ViT framework to create a hybrid object detection and tracking CP model. This extension integrates V2X-ViT’s object detection features with historical object tracking information. Object tracking features are extracted from previously tracked objects using a spatial-temporal pyramidal 3D network. Subsequently, an operation concatenates the object detection and tracking features along their channel dimension to enable robust object inferences. Experimental results demonstrate that HYDRO-3D leverages longer-term historical information from object tracking to enhance object detection accuracy.
LiDAR sensors offer precise 3D depth measurements but are costly to implement at scale. On the other hand, camera-based object detection presents a more cost-effective solution with a straightforward setup for object localization. Recognizing this advantage, the work in [215] introduced a sparse vision transformer framework named CoBEVT. This framework leverages features derived from multi-agent multi-camera sensors to collaboratively generate BEV-segmented maps. Within CoBEVT, a 3D vision transformer named FuseBEVT was introduced to attentively fuse information obtained from BEV features shared among multiple agents. A pivotal component of FuseBEVT is the Fused Axial (FAX) attention module, which facilitates sparse local and global spatial interactions across various multi-agent and multi-camera perspectives. FAX integrates a global attention mechanism to capture long-range relationships and a local attention mechanism aggregating intricate regional characteristics. CoBEVT exhibited superior performance in BEV segmentation tasks by exploiting the FAX module’s effectiveness in feature fusion.
The Point Transformer [182] is a highly effective method for learning compact point-based representations from 3D point clouds. It excels at capturing non-local interactions among points and generating permutation-invariant representations, rendering it particularly adept at aggregating point clouds from multiple vehicles. In the context of COOPERNAUT [165], a point transformer-based Representation Aggregator (RA) was introduced for information fusion purposes. This RA consists of two main components: a voxel max-pooling operation and a point transformer block. The voxel max-pooling operation is employed to spatially aggregate incoming messages close by down-sampling all the points within the same voxel grid cell. Subsequently, the point transformer block is used to fuse the multi-agent perception information. Empirical results have demonstrated that COOPERNAUT significantly enhances the robustness of self-driving strategies, particularly in scenarios involving risk-sensitive traffic situations.
The attention-based models mentioned earlier have made significant strides in cooperative information fusion, primarily focusing on spatial dimensions. However, they often operate on single frames, which can lead to difficulty detecting fast-moving objects, given the sparsity of point clouds. Additionally, the shared information may suffer from asynchrony and misalignment in both spatial and temporal dimensions due to transmission delays and sensor configuration disparities. Thus, it becomes imperative to concurrently address spatial and temporal aspects of shared information to achieve accurate perception.

### v4

@@ IV-D2 Alignment for Pose Errors
TABLE IV: Summary of the Works of Cooperative Information Alignment for Pose Errors.
@@
RANSAC-based [188] 2022 V2V LiDAR COMAP Late Detection, Location $\times$
V2VNetrobust [189] 2021 V2V LiDAR V2V-Sim Intermediate Detection, Motion Prediction $\times$
@@
RoCo [191] 2024 V2X LiDAR DAIR-V2X, V2XSet Intermediate Detection $\surd$
@@
V2X-PC [136] 2024 V2X LiDAR DAIR-V2X, V2XSet Early Detection $\times$
CoBEVGlue [193] 2024 V2X LiDAR OPV2V,DAIR-V2X,V2V4Real Intermediate Detection $\surd$
@@
FreeAlign [195] 2024 V2X LiDAR OPV2V, DAIR-V2X Intermediate Detection $\surd$
@@
V2X-ViT [145] 2022 V2X LiDAR V2XSet Intermediate Detection $\surd$
EMIFF [149] 2024 V2X Camera DAIR-V2X Intermediate Detection $\times$
CBM [197] 2024 V2V Camera,LiDAR SIND, OPV2V Late Detection $\surd$
KM-based [198] 2024 V2V Camera KITTI, PreScan Late Detection $\times$
@@
Many efforts have been undertaken to address pose errors, as summarized in Table IV. A widely adopted approach involves statistical learning models. For instance, the method presented in [188] corrected relative locations between two CAVs by matching key points in their perception results with the Random Sample Consensus (RANSAC) algorithm [200]. While effective, this approach required the setting of problem-specific thresholds. Based on V2VNet [140], a consistency module [189] was introduced to enhance relative pose estimations by structuring the global coherence of absolute poses as a Markov random field and incorporating Bayesian reweighting. Unfortunately, this model relies on ground-truth poses during the training phase, which may be limited in practical scenarios. CoAlign [190] proposed an Agent-Object Pose Graph Optimization method to improve pose consistency using a pose graph and optimization techniques. RoCo [191] modelled the pose correction problem as an object-matching task that reliably associates common objects detected by different agents, iteratively adjusting the agent poses to minimize alignment errors of the associated objects. These two methods effectively mitigated relative pose errors among agents without requiring precise supervision during training.
Point cloud registration methods typically focus on refining the Iterative Closest Point (ICP) algorithm [201], widely used for aligning 3D point clouds. In [192], an ICP-based matching algorithm aligned two point sets, one representing 3D bounding boxes detected by the Ego and the other by the sender, with the relative transformation formulated as an optimal transport problem [202] to minimize transportation costs between source and target points. V2X-PC [136] aligned point clusters from spatial and temporal dimensions and proposed parameter-free solutions that can adapt to different noisy levels without finetuning. CoBEVGlue [193] introduced a spatial alignment module, BEVGlue, which estimates relative poses between agents by matching co-visible objects across agents without relying on an external localization system. Unlike traditional point cloud registration methods, BEVGlue only requires the transmission of bounding boxes and tracking IDs, reducing communication bandwidth while maintaining high-quality matching by checking the spatial relationship between matched nodes.
@@
High-definition maps represent a valuable resource for achieving precise spatial alignment, primarily due to their ability to provide accurate self-vehicle localization. In this regard, a map-based CP framework called the “map container” was introduced in [194], which establishes benchmarks for spatiotemporal transformations by automatically mapping multi-agent perceptual information into the map coordinate space using a map-matching algorithm. Nevertheless, this algorithm relies on the support of monocular cameras and GNSS/IMU devices. FreeAlign [195] leveraged the invariant geometric structure composed of objects commonly perceived and observed by the individual agents to estimate the relative poses without relying on expensive high-end global localization.
Multi-scale feature interaction can effectively reduce feature map misalignment caused by pose errors by encoding the feature into multiple scales and then aggregating them into a final collaboration feature. In this context, a spatiotemporal awareness CP framework SCOPE [196] was designed to generate robust representations while accounting for localization errors through confidence-aware multi-scale feature interaction. Similarly, the Multi-scale Window Attention Module in V2X-ViT [145] was proposed to capture long-range spatial interactions at multiple scales using a pyramid of resolution windows, allowing for local and global spatial feature interactions and improving detection robustness against positioning errors. In EMIFF [149], the Multi-scale Cross Attention and Camera-aware Channel Masking modules are developed to address the negative effects of pose errors by selectively integrating features, applying spatial offsets, and reweighting features based on camera parameters.
Inter-agent object association aims to mitigate localization errors in pose transformation by matching key points or objects from connected agents and determining the optimal relative pose transformation. To facilitate this, the Context-based Matching (CBM) algorithm [197] generated object clusters based on bounding box headings and established correspondences through iterative pairwise comparison to maximize global consensus. The work in [198] handled spatial differences using a data association and spatial error calibration method based on the Kuhn-Munkres algorithm. However, identifying correspondences can be challenging due to visual folding, non-visible objects, and noisy perception. To address this, [199] framed correspondence identification as a graph-matching problem,
TABLE V: Summary of the Works of Cooperative Information Fusion.
@@
SiCP [206] 2024 V2X LiDAR OPV2V Intermediate Detection $\times$
CoBEVFusion [207] 2023 V2V LiDAR & Camera OPV2V Intermediate Detection, Segmentation $\times$
V2X-BGN [208] 2024 V2X Camera V2XSet Late Detection $\times$
@@
CPV-RCNN [211] 2023 V2X LiDAR CARLA Late Detection $\surd$
V2X-PC [136] 2024 V2X LiDAR DAIR-V2X, V2XSet Early Detection $\times$
@@
QUEST [215] 2024 V2I Camera DAIR-V2X Intermediate Detection $\times$
@@
V2VFormer [216] 2024 V2V LiDAR & Camera OPV2V, V2XSim, V2V4Real Intermediate Detection $\times$
V2VFormer++ [217] 2024 V2V LiDAR & Camera OPV2V, V2XSim Intermediate Detection $\times$
MKD-Cooper [218] 2023 V2V LiDAR OPV2V, V2XSim, Experiment Intermediate Detection $\surd$
V2X-AHD [219] 2023 V2X LiDAR V2Xset Intermediate Detection $\surd$
@@
Coop3D-Infra [220] 2022 V2I LiDAR CARLA Hybrid Detection $\times$
@@
Hybrid CP [222] 2023 V2V LiDAR OPV2V Hybrid Detection $\times$
employing a masked GNN to manage non-visible objects, explicitly addressing occlusion and limited FoV challenges.
IV-E Cooperative Information Fusion
@@
The effectiveness of CP crucially depends on the design of an efficient fusion strategy, which aims to consolidate shared information to provide coherent and accurate perception results. The following will examine recent fusion methods employed in CP information. For reference, Table V provides a summary of all discussed fusion models in this subsection.
IV-E1 Traditional Fusion
Primary reduction operators, such as summation, maximum, and average pooling, are widely used in cooperative information fusion for multi-agent perception. For example, the Cooper [97] was developed for CAVs to enhance 3D object detection using raw point cloud data, though it involved high data overhead. F-Cooper [203] extended Cooper to the feature level, balancing bandwidth constraints and detection accuracy by using max pooling for V2V voxel feature fusion. Similarly, PillarGrid [204] applied grid-wise max-pooling to integrate deep features for cooperative 3D object detection. CoCa3D [205] used point-wise maximum fusion to combine BEV features from multiple CAVs, enhancing camera-based detection through cooperative depth estimation. Besides using these commonly accepted reduction operators, dedicated fusion modules were designed to fuse BEV features from CAVs as well. SiCP [206] proposed a novel Dual-Perception Network (DP-Net) to assign different weights for LiDAR point cloud provided by different CAVs by extracting crucial gradient information requiring only a few parameters. CoBEVFusion [207] extended the idea by employing a dedicated fusion module to combine heterogeneous features gathered from the camera and LiDAR on two CAVs.
As the late fusion strategy requires a minimum volume of exchanged information between CAVs [211], several traditional fusion approaches employing the late fusion strategy have been explored in practice. In the context of object detection, non-maximum suppression (NMS) serves as a non-parametric method for filtering overlapping bounding boxes based on confidence scores without requiring training or fitting parameters. V2X-BGN [208] used NMS to aggregate redundant detection results from intelligent agents globally, based on 2D IoU in BEV space. A rule-based fusion scheme for CPMs was proposed in [209] to associate locally detected objects with those identified via V2X communication, designed for scenarios where CAVs generate CPMs based on ETSI standards. This scheme utilized a fusion centre to merge and broadcast CPMs to nearby CAVs for further perception tasks.
@@ IV-E1 Traditional Fusion
The non-parametric fusion methods offer minimal computational costs but often fail to capture data correlations. In contrast, parametric methods generally provide more accurate and reliable fusion by incorporating measurements and their uncertainties. For instance, the Covariance Intersection (CI) algorithm can consistently fuse estimates without prior knowledge of cross-correlations. Building on this concept, [210] introduced a CI-based fusion strategy for CP, achieving consistent fusion of local sensor data and object information from other agents through V2X communication.
In contrast to object-level and BEV-based approaches that often lead to feature degradation, the Point Cluster Aggregation (PCA) module in V2X-PC [136] preserves the low-level structural details of point clusters, thereby improving the accuracy of proposal bounding boxes. PCA integrates point clusters from multiple agents by matching them based on the distances between cluster centres and merging clusters that represent the same object into a new cluster. Its computational complexity is determined solely by the number of point clusters, without requiring convolution or padding operations, which makes it both effective and efficient for long-range collaborative perception.
IV-E2 Graph-based Fusion
@@
Multi-agent collaboration can be represented as a collaboration graph, with nodes as agent states and edges as pairwise collaborations. At the core is the GNN, which extracts relational data and generates node embeddings using local attributes and neighbouring information. Node states are continuously updated through inter-node communication and aggregation.
@@
In V2VNet [140], each vehicle used a fully connected GNN [223] as an aggregation module for data fusion from various viewpoints. Messages were aggregated at each node using a mask-aware accumulation operator, considering overlapping FoVs. However, it employed a scalar-valued edge weight to indicate the importance of the neural message from one agent to another, which could not effectively model the strength of collaboration between two agents. On the other hand, DiscoNet [144] employed an edge encoder to establish correlations between feature maps from different agents, generating matrix-valued edge weights. The edge encoder concatenated feature maps and applied convolutional layers for downsampling, followed by a softmax operation for normalization. Knowledge distillation enhanced training by aligning intermediate fusion-based student model features with the early fusion-based teacher model, improving feature abstraction and object detection performance.
The Heterogeneous Multi-agent Self-attention (HMSA) module in V2X-ViT [145] fuse cooperative feature by associating agent types with nodes and edges in a directed graph and computing importance weights between nodes accordingly. To further enhance CP performance, HYDRO-3D [212] extends the V2X-ViT to create a hybrid object detection and tracking CP model by combining V2X-ViT’s detection features with historical tracking data extracted using a spatial-temporal pyramidal 3D network. The combined features improve robustness and enhance detection accuracy, as demonstrated in experiments.
Traditional feature fusion methods, such as concatenation or summation, provide only fixed linear fusion and fail to account for the varying importance of feature maps. To address this, [213] proposed a graph attention feature fusion network that selectively emphasizes important regions within feature maps by employing both channel and spatial attention-based aggregation. This approach adaptively highlights informative areas, enhancing the ego vehicle’s feature representation and improving object detection precision.
IV-E3 Attention-based Fusion
The attention mechanism introduced in the transformer model has found widespread use in determining fusion weights. The Point Transformer [162] effectively learns compact representations from 3D point clouds, excelling at capturing non-local interactions and generating permutation-invariant representations. It is well-suited for aggregating LiDAR point clouds from multiple agents. In COOPERNAUT [146], a point transformer-based representation aggregator was introduced for information fusion, consisting of a voxel max-pooling operation for spatial aggregation and a point transformer block for fusing multi-agent perception data. Similarly, V2VFormer [216] and V2VFormer++ [217] also utilized channel and spatial information based attentions to fuse the features exacted from multiple CAVs. Together with knowledge distillation technology, several attention modules were employed to effectively transfer valuable knowledge from multiple excellent teacher models in different views and then aggregated into the student model on the ego CAV [219, 218].
Given the high cost of LiDAR sensors, the study in [214] proposed a sparse vision transformer framework, CoBEVT, which uses multi-agent multi-camera features for collaborative BEV segmentation. CoBEVT introduced a 3D vision transformer that fuses BEV features from multiple agents using the Fused Axial attention module, which integrates global and local attention mechanisms to capture both long-range and detailed regional interactions. QUEST [215] followed the variable query concept first proposed by DETR [224] and extended it to V2I CP settings for camera-based 3D object detection.